# Generalized Homogeneous Derivations on Graded Rings

### v5

@@ 4.2. Results on generalized homogeneous derivations
Assume $F=0$. For any elements $x,y\in R$, we have $F(xy)=0$. Since
$F(xy)=F(x)y+xd(y),$
it follows that $xd(y)=0$ for all $x,y\in R$. Hence $xRd(y)=\{0\}$ for all $x,y\in R$. In particular, for some nonzero homogeneous element $r\in\mathcal{H}(R)\setminus\{0\}$, we have $rRd(y)=\{0\}$ for all $y\in R$. According to [1, Proposition 2.1], this implies $d(y)=0$ for all $y\in R$. Thus $d=0$, which contradicts the assumption. ∎
In [2], it was shown that a prime ring $R$ with a nonzero ideal $\mathfrak{I}$ is commutative if it admits a generalized derivation $F$ satisfying
$F(xy)\pm xy\in Z(R)\quad\text{or}\quad F(x)F(y)\pm xy\in Z(R)$
for all $x,y\in\mathfrak{I}$. We now extend this result to gr-prime rings in the context of generalized homogeneous derivations.
@@
Consider the case
$F(xy)-xy\in Z(R)$
for all $x,y\in\mathfrak{I}$. Using the same reasoning as in the proof of [2, Theorem 2.1], we obtain the identity
$z,z_{1}]xyd(z)=0$
for all $x,y,z,z_{1}\in\mathfrak{I}$, which yields $z,z_{1}]xR\mathfrak{I}d(z)=\{0\}$ for all $x,z,z_{1}\in\mathfrak{I}$. Since $\mathfrak{I}$ is a graded ideal of $R$, we have
$r^{\prime},z_{1}]xRrd(r^{\prime})=\{0\}$
for all $x,z_{1}\in\mathfrak{I}$ and $r,r^{\prime}\in\mathfrak{I}\cap\mathcal{H}(R)$. According to [1, Proposition 2.1], either $z_{1},r^{\prime}]x=0$ or $rd(r^{\prime})=0$ for all $x,z_{1}\in\mathfrak{I}$ and $r,r^{\prime}\in\mathfrak{I}\cap\mathcal{H}(R)$. This implies that
$z,z_{1}]\mathfrak{I}=\{0\}\quad\text{or}\quad\mathfrak{I}d(z)=\{0\}$
for all $z,z_{1}\in\mathfrak{I}$. Define
@@
Then $\mathfrak{I}_{1}$ and $\mathfrak{I}_{2}$ are additive subgroups of $\mathfrak{I}$ with $\mathfrak{I}=\mathfrak{I}_{1}\cup\mathfrak{I}_{2}$. Since a group cannot be expressed as the union of two proper subgroups, either $\mathfrak{I}_{1}=\mathfrak{I}$ or $\mathfrak{I}_{2}=\mathfrak{I}$. We consider these cases separately.
Case 1: If $z_{1},z]\mathfrak{I}=\{0\}$ for all $z,z_{1}\in\mathfrak{I}$, then, since $\mathfrak{I}$ is an ideal, we obtain $z_{1},z]R\mathfrak{I}=\{0\}$ for all $z,z_{1}\in\mathfrak{I}$. As $\mathfrak{I}$ is a nonzero graded ideal, there exists $r\in\mathfrak{I}\cap\mathcal{H}(R)\setminus\{0\}$ such that
$z,z_{1}]Rr=\{0\}$
for all $z,z_{1}\in\mathfrak{I}$. By [1, Proposition 2.1], it follows that $z,z_{1}]=0$ for all $z,z_{1}\in\mathfrak{I}$. Hence $\mathfrak{I}$ is commutative. Therefore, $R$ is commutative.
Case 2: If $\mathfrak{I}d(z)=\{0\}$ for all $z\in\mathfrak{I}$, then $\mathfrak{I}Rd(z)=\{0\}$ for all $z\in\mathfrak{I}$. In particular, $rRd(z)=\{0\}$ for some $r\in\mathfrak{I}\cap\mathcal{H}(R)\setminus\{0\}$ and all $z\in\mathfrak{I}$. Using [1, Proposition 2.1], we obtain $d(z)=0$ for all $z\in\mathfrak{I}$. Hence $d$ vanishes on $\mathfrak{I}$. By [1, Lemma 2.6], $d$ is zero on $R$, which is a contradiction.
For the second case $F(xy)+xy\in Z(R)$ for all $x,y\in\mathfrak{I}$, the argument reduces to the first case by considering F-F instead of $F$. ∎
Next, we extend [2, Theorem 2.5] to gr-prime rings by considering a pair of generalized homogeneous derivations $F_{1}$ and $F_{2}$ satisfying
$F_{1}(x)F_{2}(y)\pm xy\in Z(R)$
for all $x,y$ in a graded ideal $\mathfrak{I}$ of $R$.
@@
(7) $F_{1}(x)F_{2}(y)-xy\in Z(R)$
for all $x,y\in\mathfrak{I}$. Substituting $yz$ for $y$ in (7), we obtain
(8) $\big(F_{1}(x)F_{2}(y)-xy\big)z+F_{1}(x)yd_{2}(z)\in Z(R)$
for all $x,y\in\mathfrak{I}$ and $z\in R$. Taking the commutator of (8) with $z$, we obtain
(9) $F_{1}(x)[yd_{2}(z),z]+[F_{1}(x),z]yd_{2}(z)=0$
for all $x,y\in\mathfrak{I}$ et $z\in R$. Substituting $F_{1}(x)y$ for $y$ in (9), we arrive at
(10) $F_{1}(x),z]F_{1}(x)yd_{2}(z)=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. This implies $F_{1}(x),z]F_{1}(x)R\mathfrak{I}d_{2}(z)=\{0\}$ for all $x\in\mathfrak{I}$ et $z\in R$. In particular,
$F_{1}(r),r^{\prime}]F_{1}(r)R\mathfrak{I}d_{2}(r^{\prime})=\{0\}$
for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and $r^{\prime}\in\mathcal{H}(R)$. According to [1, Proposition 2.1], either
$F_{1}(r),r^{\prime}]F_{1}(r)=0\quad\text{or}\quad\mathfrak{I}d_{2}(r^{\prime})=\{0\}$
for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and $r^{\prime}\in\mathcal{H}(R)$. Thus, $F_{1}(x),z]F_{1}(x)=0$ or $\mathfrak{I}d_{2}(z)=\{0\}$ for all $x\in\mathfrak{I}$ and $z\in R$. Let
@@
Clearly, $\mathfrak{J}_{1}$ and $\mathfrak{J}_{2}$ are additive subgroups of $R$ whose union is $R$. Since a group cannot be the union of two proper subgroups, either $\mathfrak{J}_{1}=R$ or $\mathfrak{J}_{2}=R$.
If $\mathfrak{J}_{2}=R$, then $\mathfrak{I}d_{2}(z)=\{0\}$ for all $z\in R$. Since $\mathfrak{I}$ is an ideal, $\mathfrak{I}Rd_{2}(z)=\{0\}$ for all $z\in R$. In particular, $rRd_{2}(z)=\{0\}$ for all $z\in R$ and some $r\in\mathfrak{I}\cap\mathcal{H}(R)\setminus\{0\}$. According to [1, Proposition 2.1], we conclude that $d_{2}(z)=0$ for all $z\in R$. Hence $d_{2}=0$, which contradicts our assumption. Therefore, $F_{1}(x),z]F_{1}(x)=0$ for all $x\in\mathfrak{I}$ and $z\in R$. Replacing $z$ by $zz^{\prime}$, we obtain $F_{1}(x),z]z^{\prime}F_{1}(x)=0$ for all $x\in\mathfrak{I}$ and $z,z^{\prime}\in R$, which implies $F_{1}(x),z]RF_{1}(x)=\{0\}$ pour tous $x\in\mathfrak{I}$ et $z\in R$. In particular,
$F_{1}(r),z]RF_{1}(r)=\{0\}$
for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and $z\in R$. Invoking [1, Proposition 2.1], we conclude that either $F_{1}(r)=0$ or $F_{1}(r),z]=0$ for all $z\in R$. Hence $F_{1}(x)=0$ or $F_{1}(x),z]=0$ for all $x\in\mathfrak{I}$ and $z\in R$. In both cases, we obtain $F_{1}(x),z]=0$ for all $x\in\mathfrak{I}$ and $z\in R$. Replacing $x$ by $xz$, we obtain
(11) $x[d_{1}(z),z]+[x,z]d_{1}(z)=0$
for all $x\in\mathfrak{I}$ and $z\in R$. Substituting $sx$ for $x$ in (11), we arrive at $s,z]xd_{1}(z)=0$ for all $x\in\mathfrak{I}$ and $s,z\in R$, which implies $s,z]R\mathfrak{I}d_{1}(z)=\{0\}$ for all $s,z\in R$. Using similar arguments as above, either $[s,z]=0$ or $d_{1}(z)=0$ for all $s,z\in R$. Since $d_{1}\neq 0$, we must have $[s,z]=0$ for all $s,z\in R$. Therefore, $R$ is commutative.
For the second case $F_{1}(x)F_{2}(y)+xy\in Z(R)$ for all $x,y\in\mathfrak{I}$, the argument reduces to the first case by considering $F_{1}$ instead of $F_{1}$. ∎
The following example shows that the gr-primeness hypothesis cannot be omitted from the above theorems.
@@
Let $R=\mathbb{C}[t_{1},t_{2},t_{3}]\times\left\{\left(\begin{array}[]{cc}a&b\\ 0&0\end{array}\right)\Big|\;a,b\in\mathbb{C}\right\}$ with $\mathbb{Z}\times\mathbb{Z}_{2}$-grading. Then $R$ is not gr-prime. Let $\mathfrak{I}=\mathbb{C}[t_{1},t_{2},t_{3}]\times\left\{\left(\begin{array}[]{cc}0&a\\ 0&0\end{array}\right)\;\Big|\;a\in\mathbb{C}\right\}.$ Clearly, $\mathfrak{I}$ is a nonzero graded ideal of $R$. Consider the mappings:
$\begin{array}[]{cccc}F_{1}:&R&\rightarrow&R\\ &(f,M)&\mapsto&\big(t_{3}\big(f+\frac{\partial f}{\partial t_{3}}\big),0\big)\end{array},\qquad\begin{array}[]{cccc}F_{2}=d_{2}:&R&\rightarrow&R\\ &(f,M)&\mapsto&\big(t_{1}\frac{\partial f}{\partial t_{2}},0\big)\end{array},$
and
$\begin{array}[]{cccc}d_{1}:&R&\rightarrow&R\\ &(f,M)&\longmapsto&\big(t_{2}t_{3}\frac{\partial f}{\partial t_{3}},0\big).\end{array}$
Then $(F_{1},d_{1})_{h}$ and $(F_{2},d_{2})_{h}$ are generalized homogeneous derivations on $R$. Moreover, $F_{1}(xy)\pm xy\in Z(R)$ and $F_{1}(x)F_{2}(y)\pm xy\in Z(R)$ for all $x,y\in\mathfrak{I}$. However, $R$ is noncommutative.
@@ 5. Existence Conditions for Central Graded Ideals in Gr-Semiprime Rings
In this section, we investigate the behavior of graded ideals under generalized homogeneous derivations, and we characterize when such rings necessarily contain nonzero central graded ideals.
In [3], it was shown that if a ring $R$ admits generalized derivations $F_{1}$ and $F_{2}$ with associated nonzero derivations $d_{1}$ and $d_{2}$, respectively, such that
$F_{1}(x)x\pm xF_{2}(x)=0$
for all $x\in\mathfrak{I}$, where $\mathfrak{I}$ is a nonzero ideal of $R$, then $R$ contains a nonzero central ideal. We extend this result to the graded case by studying generalized homogeneous derivations $F_{1}$ and $F_{2}$ satisfying
$F_{1}(x)y\pm xF_{2}(y)\in Z(R)$
for all $x,y\in\mathfrak{I}$, where $\mathfrak{I}$ is a graded ideal of a gr-semiprime ring $R$.
@@
(12) $F_{1}(x)y-xF_{2}(y)\in Z(R)$
for all $x,y\in\mathfrak{I}$. Substituting $yz$ for $y$ in (12), we obtain
(13) $\big(F_{1}(x)y-xF_{2}(y)\big)z-xyd_{2}(z)\in Z(R)$
for all $x,y,z\in\mathfrak{I}$. Taking the commutator of (13) with $z$ yields
(14) $xy[d_{2}(z),z]+x[y,z]d_{2}(z)+[x,z]yd_{2}(z)=0$
for all $x,y,z\in\mathfrak{I}$. Substituting $d_{2}(z)x$ for $x$ in (14), we obtain
(15) $d_{2}(z),z]xyd_{2}(z)=0$
for all $x,y,z\in\mathfrak{I}$. Substituting $d_{2}(z)y$ for $y$ in (15), we obtain
(16) $d_{2}(z),z]xd_{2}(z)yd_{2}(z)=0$
for all $x,y,z\in\mathfrak{I}$. By subtracting (16) from (15) and using the previous identities, we get
$d_{2}(z),z]x[d_{2}(z),z]y[d_{2}(z),z]=0$
for all $x,y,z\in\mathfrak{I}$. This implies
@@
for all $z\in\mathfrak{I}$. Since $R$ is gr-semiprime, there exists a family $\mathcal{F}:=\{P_{i}\mid i\in\Lambda\}$ of gr-prime ideals such that $\bigcap_{i\in\Lambda}P_{i}=\{0\}$. Therefore,
$d_{2}(z),z]\mathfrak{I}[d_{2}(z),z]\mathfrak{I}[d_{2}(z),z]\subseteq P_{i}$
for all $i\in\Lambda$ and all $z\in\mathfrak{I}$. By [1, Proposition 2.1], and since $\mathfrak{I}$ is a graded ideal, we have $d_{2}(z),z]\in P_{i}$ for all $i\in\Lambda$ and $z\in\mathfrak{I}$. Hence $d_{2}(z),z]=0$ for all $z\in\mathfrak{I}$. Thus, by [1, Theorem 4.1], $R$ contains a nonzero central graded ideal.
For the second case
$F_{1}(x)y+xF_{2}(y)\in Z(R)$
for all $x,y\in\mathfrak{I}$, the argument reduces to the first case by considering $F_{2}$ instead of $F_{2}$. ∎
@@
Using similar arguments with appropriate modifications, and considering the cases $F_{1}=F_{2}$ or $F_{1}=-F_{2}$ in Theorem 5.1, we obtain the following result. This extends the graded version of Posner’s Second Theorem [1, Theorem 3.3] to generalized homogeneous derivations on gr-prime rings, providing a characterization of commutativity.
@@
$R=\left\{\left(\begin{array}[]{cc}a&b\\ 0&0\end{array}\right)\Big|\;a,b\in\mathbb{R}\right\}$
@@
$\begin{array}[]{cccc}F_{1}:&R&\longrightarrow&R\\ &\left(\begin{array}[]{cc}a&b\\ 0&0\end{array}\right)&\longmapsto&\left(\begin{array}[]{cc}a&2b\\ 0&0\end{array}\right)\end{array}\quad\text{and}\quad\begin{array}[]{cccc}F_{2}:&R&\longrightarrow&R\\ &\left(\begin{array}[]{cc}a&b\\ 0&0\end{array}\right)&\longmapsto&\left(\begin{array}[]{cc}0&a+2b\\ 0&0\end{array}\right)\end{array}$
@@
$d_{1}\left(\begin{array}[]{cc}a&b\\ 0&0\end{array}\right)=\left(\begin{array}[]{cc}0&b\\ 0&0\end{array}\right),\qquad d_{2}=F_{2}.$
Let
$\mathfrak{I}=\left\{\left(\begin{array}[]{cc}0&a\\ 0&0\end{array}\right)\;\Big|\;a\in\mathbb{R}\right\}$
be a graded ideal of $R$. Even though $F_{1}$ and $F_{2}$ satisfy the conditions of Theorem 5.1, the ring $R$ has no nonzero central graded ideal.

### v6

@@ 4.2. Results on generalized homogeneous derivations
Assume $F=0$. For any elements $x,y\in R$, the relation $F(xy)=F(x)y+xd(y)$ yields $xd(y)=0$ for all $x,y\in R$. Hence $xRd(y)=\{0\}$ for all $x,y\in R$. For some nonzero $r\in\mathcal{H}(R)\setminus\{0\}$, we have $rRd(y)=\{0\}$ for all $y\in R$. By [1, Proposition 2.1], $d(y)=0$ for all $y\in R$, contradicting $d\neq 0$. ∎
In [2], it was shown that a prime ring $R$ with a nonzero ideal $\mathfrak{I}$ is commutative if it admits a generalized derivation $F$ satisfying $F(xy)\pm xy\in Z(R)$ or $F(x)F(y)\pm xy\in Z(R)$ for all $x,y\in\mathfrak{I}$. We now extend this result to gr-prime rings in the context of generalized homogeneous derivations.
@@
Consider the case $F(xy)-xy\in Z(R)$ for all $x,y\in\mathfrak{I}$. Using the same reasoning as in the proof of [2, Theorem 2.1], we obtain the identity $z,z_{1}]xyd(z)=0$ for all $x,y,z,z_{1}\in\mathfrak{I}$, which yields $z,z_{1}]xR\mathfrak{I}d(z)=\{0\}$. Since $\mathfrak{I}$ is a graded ideal, we have $r^{\prime},z_{1}]xRrd(r^{\prime})=\{0\}$ for all $x,z_{1}\in\mathfrak{I}$ and $r,r^{\prime}\in\mathfrak{I}\cap\mathcal{H}(R)$. By [1, Proposition 2.1], either $z_{1},r^{\prime}]x=0$ or $rd(r^{\prime})=0$. This implies either $z,z_{1}]\mathfrak{I}=\{0\}$ or $\mathfrak{I}d(z)=\{0\}$ for all $z,z_{1}\in\mathfrak{I}$. Define
@@
Since a group cannot be expressed as the union of two proper subgroups, either $\mathfrak{I}_{1}=\mathfrak{I}$ or $\mathfrak{I}_{2}=\mathfrak{I}$.
Case 1. If $z_{1},z]\mathfrak{I}=\{0\}$ for all $z,z_{1}\in\mathfrak{I}$, then $z_{1},z]R\mathfrak{I}=\{0\}$. There exists $r\in\mathfrak{I}\cap\mathcal{H}(R)\setminus\{0\}$ such that $z,z_{1}]Rr=\{0\}$. By [1, Proposition 2.1], $z,z_{1}]=0$ for all $z,z_{1}\in\mathfrak{I}$, so $\mathfrak{I}$ is commutative and $R$ is commutative.
Case 2. If $\mathfrak{I}d(z)=\{0\}$ for all $z\in\mathfrak{I}$, then $\mathfrak{I}Rd(z)=\{0\}$. For some $r\in\mathfrak{I}\cap\mathcal{H}(R)\setminus\{0\}$, $rRd(z)=\{0\}$ for all $z\in\mathfrak{I}$. By [1, Proposition 2.1], $d(z)=0$ for all $z\in\mathfrak{I}$, hence by [1, Lemma 2.6] $d=0$ on $R$, a contradiction.
For the case $F(xy)+xy\in Z(R)$, the argument reduces to the first case by considering F-F instead of $F$. ∎
@@
$F_{1}(x)F_{2}(y)-xy\in Z(R)$ (4.6)
for all $x,y\in\mathfrak{I}$. Substituting $yz$ for $y$ in (4.6):
$\bigl(F_{1}(x)F_{2}(y)-xy\bigr)z+F_{1}(x)yd_{2}(z)\in Z(R)$ (4.7)
for all $x,y\in\mathfrak{I}$ and $z\in R$. Taking the commutator of (4.7) with $z$:
$F_{1}(x)[yd_{2}(z),z]+[F_{1}(x),z]yd_{2}(z)=0.$ (4.8)
Substituting $F_{1}(x)y$ for $y$ in (4.8):
$F_{1}(x),z]F_{1}(x)yd_{2}(z)=0$ (4.9)
for all $x,y\in\mathfrak{I}$ and $z\in R$. This implies $F_{1}(r),r^{\prime}]F_{1}(r)R\mathfrak{I}d_{2}(r^{\prime})=\{0\}$ for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and $r^{\prime}\in\mathcal{H}(R)$. By [1, Proposition 2.1], either $F_{1}(x),z]F_{1}(x)=0$ or $\mathfrak{I}d_{2}(z)=\{0\}$ for all $x\in\mathfrak{I}$ and $z\in R$. Let
@@
Since $\mathfrak{J}_{1}\cup\mathfrak{J}_{2}=R$ and a group cannot be the union of two proper subgroups, either $\mathfrak{J}_{1}=R$ or $\mathfrak{J}_{2}=R$.
If $\mathfrak{J}_{2}=R$, arguing as in the proof of Theorem 4.6, we conclude $d_{2}=0$, a contradiction. Therefore $F_{1}(x),z]F_{1}(x)=0$ for all $x\in\mathfrak{I}$ and $z\in R$. Replacing $z$ by $zz^{\prime}$, we get $F_{1}(x),z]RF_{1}(x)=\{0\}$, so by [1, Proposition 2.1], $F_{1}(r)=0$ or $F_{1}(r),z]=0$ for all $z\in R$. In both cases $F_{1}(x),z]=0$. Replacing $x$ by $xz$:
$x[d_{1}(z),z]+[x,z]d_{1}(z)=0$ (4.10)
for all $x\in\mathfrak{I}$ and $z\in R$. Substituting $sx$ for $x$: $s,z]xd_{1}(z)=0$, so $s,z]R\mathfrak{I}d_{1}(z)=\{0\}$. Either $[s,z]=0$ or $d_{1}(z)=0$ for all $s,z\in R$. Since $d_{1}\neq 0$, we have $[s,z]=0$ for all $s,z\in R$, so $R$ is commutative. The case $F_{1}(x)F_{2}(y)+xy\in Z(R)$ reduces to the above by considering $F_{1}$ instead of $F_{1}$. ∎
The following example shows that the gr-primeness hypothesis cannot be omitted.
@@
Let
$R=\mathbb{C}[t_{1},t_{2},t_{3}]\times\left\{\begin{pmatrix}a&b\\ 0&0\end{pmatrix}\;\Big|\;a,b\in\mathbb{C}\right\}$
with $\mathbb{Z}\times\mathbb{Z}_{2}$-grading. Then $R$ is not gr-prime. Let $\mathfrak{I}=\mathbb{C}[t_{1},t_{2},t_{3}]\times\left\{\begin{pmatrix}0&a\\ 0&0\end{pmatrix}\;\Big|\;a\in\mathbb{C}\right\}$. Clearly, $\mathfrak{I}$ is a nonzero graded ideal of $R$. Consider the mappings:
$F_{1}(f,M)=\Bigl(t_{3}\Bigl(f+\frac{\partial f}{\partial t_{3}}\Bigr),0\Bigr),\qquad F_{2}(f,M)=d_{2}(f,M)=\Bigl(t_{1}\frac{\partial f}{\partial t_{2}},0\Bigr),\qquad d_{1}(f,M)=\Bigl(t_{2}t_{3}\frac{\partial f}{\partial t_{3}},0\Bigr).$
Then $(F_{1},d_{1})_{h}$ and $(F_{2},d_{2})_{h}$ are generalized homogeneous derivations on $R$ satisfying $F_{1}(xy)\pm xy\in Z(R)$ and $F_{1}(x)F_{2}(y)\pm xy\in Z(R)$ for all $x,y\in\mathfrak{I}$. However, $R$ is noncommutative.
@@ 5. Existence conditions for central graded ideals in gr-semiprime rings
In this section, we investigate the behavior of graded ideals under generalized homogeneous derivations and characterize when such rings necessarily contain nonzero central graded ideals.
In [3], it was shown that if a ring $R$ admits generalized derivations $F_{1}$ and $F_{2}$ with associated nonzero derivations $d_{1}$ and $d_{2}$, respectively, such that $F_{1}(x)x\pm xF_{2}(x)=0$ for all $x\in\mathfrak{I}$, where $\mathfrak{I}$ is a nonzero ideal of $R$, then $R$ contains a nonzero central ideal. We extend this result to the graded case.
@@
$F_{1}(x)y-xF_{2}(y)\in Z(R)$ (5.1)
for all $x,y\in\mathfrak{I}$. Substituting $yz$ for $y$ in (5.1):
$\bigl(F_{1}(x)y-xF_{2}(y)\bigr)z-xyd_{2}(z)\in Z(R)$ (5.2)
for all $x,y,z\in\mathfrak{I}$. Taking the commutator of (5.2) with $z$:
$xy[d_{2}(z),z]+x[y,z]d_{2}(z)+[x,z]yd_{2}(z)=0.$ (5.3)
Substituting $d_{2}(z)x$ for $x$ in (5.3):
$d_{2}(z),z]xyd_{2}(z)=0.$ (5.4)
Substituting $d_{2}(z)y$ for $y$ in (5.4):
$d_{2}(z),z]xd_{2}(z)yd_{2}(z)=0.$ (5.5)
Subtracting (5.5) from (5.4) and using the previous identities:
@@
for all $z\in\mathfrak{I}$. Since $R$ is gr-semiprime, there exists a family $\mathcal{F}:=\{P_{i}\mid i\in\Lambda\}$ of gr-prime ideals with $\bigcap_{i\in\Lambda}P_{i}=\{0\}$. By [1, Proposition 2.1], $d_{2}(z),z]\in P_{i}$ for all $i\in\Lambda$ and $z\in\mathfrak{I}$, hence $d_{2}(z),z]=0$ for all $z\in\mathfrak{I}$. By [1, Theorem 4.1], $R$ contains a nonzero central graded ideal. The case $F_{1}(x)y+xF_{2}(y)\in Z(R)$ reduces to the first by considering $F_{2}$ instead of $F_{2}$. ∎
@@
$R=\left\{\begin{pmatrix}a&b\\ 0&0\end{pmatrix}\;\Big|\;a,b\in\mathbb{R}\right\}$
@@
$F_{1}\begin{pmatrix}a&b\\ 0&0\end{pmatrix}=\begin{pmatrix}a&2b\\ 0&0\end{pmatrix},\qquad F_{2}\begin{pmatrix}a&b\\ 0&0\end{pmatrix}=\begin{pmatrix}0&a+2b\\ 0&0\end{pmatrix}$,
@@
$d_{1}\begin{pmatrix}a&b\\ 0&0\end{pmatrix}=\begin{pmatrix}0&b\\ 0&0\end{pmatrix},\qquad d_{2}=F_{2}.$
Let $\mathfrak{I}=\left\{\begin{pmatrix}0&a\\ 0&0\end{pmatrix}\;\Big|\;a\in\mathbb{R}\right\}$ be a graded ideal of $R$. Even though $F_{1}$ and $F_{2}$ satisfy the conditions of Theorem 5.1, the ring $R$ has no nonzero central graded ideal.